How many liters of each of a 10% acid solution and a 25% acid solution must be used to produce 60 liters of a 15% acid solution?
Answer by stanbon(75887)

How many liters of each of a 10% acid solution and a 25% acid solution must be used to produce 60 liters of a 15% acid solution?
--------------
Equation:
acid + acid = acid
0.10x + 0.25(60-x) = 0.15*60
------
Multiply thru by 100 to get:
10x + 25*60-25x = 15*60
-15x = -10*50
x = (2/3)*50
x = 33 1/3 liters (amt. of 10% solution needed)
-----------
60-x = 26 2/3 liters (amt. of 25% solution needed)
=====================================================
Cheers,
Stan H.
============